What concentration of hydroxide ions exist in a 0.008 F solution of NH3 at 25°C .


1
Expert's answer
2021-03-15T09:06:45-0400

In solution, NH3 dissociates as follows:

NH3 + H2O = NH4+ + OH-

Dissociation constant:

Kb = [NH4+][OH-] / [NH3]

From here:

[OH-] = [NH4+] = (Kb × [NH3])1/2 = (1.8 × 10-5 × 0.008 F) = 3.8 × 10-4 F


Answer: 3.8 × 10-4 F